# Dynamic EventNeRF: Reconstructing General Dynamic Scenes from Multi-view RGB and Event Streams

### v3

@@ Abstract
Volumetric reconstruction of dynamic scenes is an important problem in computer vision. It is especially challenging in poor lighting and with fast motion. This is partly due to limitations of RGB cameras: To capture frames under low lighting, the exposure time needs to be increased, which leads to more motion blur. In contrast, event cameras, which record changes in pixel brightness asynchronously, are much less dependent on lighting, making them more suitable for recording fast motion. We hence propose the first method to spatiotemporally reconstruct a scene from sparse multi-view event streams and sparse RGB frames. We train a sequence of cross-faded time-conditioned NeRF models, one per short recording segment. The individual segments are supervised with a set of event- and RGB-based losses and sparse-view regularisation. We assemble a real-world multi-view camera rig with six static event cameras around the object and record a benchmark multi-view event stream dataset of challenging motions. Our work outperforms RGB-based baselines, producing state-of-the-art results, and opens up the topic of multi-view event-based reconstruction as a new path for fast scene capture beyond RGB cameras.
@@ 1 Introduction
In this work, we explore the use of multi-view event data for event-based 4D reconstruction of non-rigid motion. Since a multi-view event camera setup is of course more complex than a monocular setup, we believe that investigating this uncharted territory can prove a helpful guideline in deciding whether the added cost is worth the gains in quality.
@@ 2 Related Works
Novel view synthesis has seen recent progress [26, 51, 23, 27, 17], with methods based on Neural Radiance Fields (NeRF) [27] being especially successful. NeRF has been adapted for dynamic scene reconstruction, e.g. head avatars [45], full-body avatars [41, 22, 12], hands [32] and other domains. Other techniques targets general scene reconstruction, which can be divided into monocular [43, 40, 52, 20] and multi-view methods [9, 4, 50, 42, 58, 54]. While the monocular methods can reconstruct the input video with high fidelity, the ambiguities of monocular input make it hard for them to learn the correct geometry, greatly limiting how far novel views can deviate from the training viewpoints. Moreover, many monocular methods greatly depend on the camera moving faster than the object, constraining the scenes they can reconstruct [11]. As we capture fast motion, moving the camera even faster is not a viable option. Our method uses a fixed camera multi-view setup. Multi-view methods generalise better to novel views, but are still limited by their RGB input, which is prone to motion blur and noise, especially with fast motion or low lighting conditions. Instead, we use event cameras, possessing both high temporal resolution and high dynamic range. To train NeRF models, one typically needs dozens of ground-truth RGB views, but in our setting, we only have six views available. Follow-up works of NeRF aim at enabling few-view reconstruction, using extensive regularisation [36], pre-training [62, 33], semantic consistency [15], or depth priors [34, 47]. Our approach (Sec. 4.3) uses a combination of regularisers and volume clipping, without a need for pretraining or additional priors, and is applicable to general scenes.
@@
EvDNeRF [1] allows to synthesise events from novel viewpoints using multi-view event streams through the deformation of a canonical volume. It does not model the appearance, needs as much as 18 views, and the deformation-based approach significantly limits the scenes and motions that can be reconstructed. In contrast, our Dynamic EventNeRF allows to reconstruct general dynamic scenes directly in RGB space while using as few as three views.
@@ 3.1 Event Camera Model and Deblurring
Instead of capturing dense RGB frames at regular intervals, event cameras operate asynchronously per each individual pixel: When a certain pixel gets brighter or darker by a certain threshold, it emits an event, which is a tuple $(t,x,y,p)$ of its timestamp $t$, pixel position $x,y$, and polarity $p\in\{-1,+1\}$. The event can be interpreted as the assertion
@@
where $\log I_{x,y}(\tau)$ is the intensity of pixel $(x,y)$ at time $\tau$ and $t^{\prime}<t$ is the timestamp of the previous event in that pixel. We call $C_{\mathrm{+1}},C_{\mathrm{-1}}>0$ the event generation thresholds. Event-based Single Integral (ESI) [49] connects the difference between logarithmic intensities on its left side and the accumulated events on the right side by summing Eq. 1 over the time. ESI can be further extended to Event-based Double Integral (EDI) [37] by averaging the intensity over the exposure period, to model motion blur caused by the traditional camera shutter. Given events $\mathbf{E}$ and the blurry image $\mathbf{B}$, one can solve a system of equations generated by EDI for the unknown instantaneous intensity image $\mathbf{I}$, thus deblurring $\mathbf{B}$. This way, EDI can be used for deblurring the RGB stream through events. To synthesise arbitrary instantaneous frames $\mathbf{I}(t)$, one could accumulate events on top of the closest recovered deblurred frame $\mathbf{I}(t_{0})$ with ESI. To improve the speed of deblurring and frame synthesis, Lin et al. propose Fast EDI [21], which uses the same event camera model and a special data structure to accelerate the queries.
@@ 4.2 Multi-Segment Model
Using one single MLP to represent longer sequences is challenging due to the capacity-performance trade-off and the ghosting artefacts caused by using a shared network. Therefore, we propose to split the full sequence into short segments and train one model per such sequence. To eliminate flickering caused by changing from one model to the next, we overlap the segments by $10\%$ at the starts and at the ends. In the overlapped parts, we render both models as $\mathbf{\hat{C}}_{i}(t)$ and $\mathbf{\hat{C}}_{i+1}(t)$ and cross-fade the predictions:
@@ 4.3 Supervision for One Segment
For sparse-view reconstruction, we use the following combination of regularisation techniques: First, we want to prevent the model from learning artefacts in the regions where only one view observes them. These areas are near and far away from the camera but still in its frustum. Suppose the views are distributed 360°360°360  ° around the object at roughly similar height: The intersection of all frustums can be approximated by a cylinder of radius $r$ that extends vertically from $y_{\mathrm{min}}$ to $y_{\mathrm{max}}$. We can then clip the reconstruction volume by setting $\sigma(x,y,z):=0$ wherever $x^{2}+z^{2}>r$, or $y>y_{\textrm{max}}$ or $y<y_{\textrm{min}}$. We also modify the ray depth sampling so that the samples are only chosen inside the cylinder. Following EventNeRF, we randomly offset the ray direction within a pixel to improve generalisation to novel views. This cylinder clipping and the improved sampling procedure result in a strong prior significantly reducing the ambiguity of the sparse-view reconstruction.
@@
Learning high-frequency details from the start is not optimal, especially in the sparse-view setting. Thus we implement coarse-to-fine training, by applying positional encoding annealing [39] both across spatial frequencies and time. Annealing improves temporal and spatial smoothness as the method starts with a coarse representation, slowly progressing towards the finer representation.
@@ 4.4 Event Accumulation
$E_{x,y}(t_{0},t_{1})=\sum_{i=i_{\mathrm{start}}}^{i_{\mathrm{end}}}\delta_{x_{i},y_{i}}^{x,y}p_{i}C_{p_{i}}b^{i_{\mathrm{end}}-i_{\mathrm{start}}-i},~{}b=0.93$ (11)
@@
In Eq. 6, we defined $\mathcal{L}_{\mathrm{event}}$, which uses $E(t_{0},t_{1})$ for random $t_{0}$ and $t_{1}$ each training iteration. As event streams could contain millions of events per second, implementing the accumulation procedure directly by processing each event on every query, would be too slow to run every training iteration. To overcome this issue, we use an approach inspired by Lin et al. [21] to accelerate the queries: We treat all pixels independently and store all intermediate accumulation results in arrays $A^{x_{i},y_{i}}_{j}=E(0,t_{i})_{x_{i},y_{i}}$, and $T^{x_{i},y_{i}}_{j}=t_{i}$, where $j$ is the per-pixel event counter. Then, once we need to compute $E(t_{0},t_{1})$, we split it into two queries:
@@ 5.5 Ablation and Design Choice Study
We conduct an ablation and design choice study (Tabs. 2, 5 and 4; per-scene in Tab. 4 in the appendix) using both synthetic and real data. In particular, we compare different options for the core model: MLP (“Full Model”), NGP [31], TensoRF-CP [7], and HexPlane [4]. NGP fails to handle the sparse-view setting. TensoRF-CP and HexPlane are both grid-based: Individual timestamps are modelled separately, so that information from one timestamp cannot propagate to other timestamps, leading to blurriness and artefacts. On the other hand, our full model with temporally-conditioned MLP explicitly controls how much information is shared between timestamps through positional encoding. Supervising one timestamp also affects others in its vicinity, resulting in better use of the training data. Disabling $\mathcal{L}_{\text{event}}$, $\mathcal{L}_{\text{acc}}$, or $\mathcal{L}_{\text{RGB}}$ leads to increased blurriness because disabling either of them will reduce the amount of supervision: Without $\mathcal{L}_{\text{event}}$, there is no way to correct relative changes between predictions through events, without $\mathcal{L}_{\text{acc}}$ there is no way to propagate sharp RGB information and static parts of the scene. And without $\mathcal{L}_{\text{RGB}}$, the model is not encouraged to predict the reference RGB frames correctly. Disabling both $\mathcal{L}_{\text{RGB}}$ and $\mathcal{L}_{\text{acc}}$ removes the entire RGB supervision. As the model only uses the events, there can be an exposure ambiguity, leading to lower metrics. Disabling both $\mathcal{L}_{\text{RGB}}$ and $\mathcal{L}_{\text{event}}$ leaves only $\mathcal{L}_{\text{acc}}$. Despite using both RGB and events, this baseline performs worse than the full model, presumably due to accumulation artefacts. Cylinder clipping prevents the model from creating geometry in regions that are only supervised by one view. If disabled, the model produces numerous artefacts in the novel views; in the case of real data specifically, it even fails to converge.$\mathcal{L}_{\text{sparsity}}$ forces the model to reduce the number of opaque pixels in the reconstruction Removing it leads to semi-transparent floating artefacts in the reconstructed volume. Using only one single model for the entire sequence duration (as opposed to dividing the sequence into shorter chunks) results in blurrier reconstruction. In this case, artefacts produced by reconstructing one of the segments can propagate to other segments, which is not the case in the multi-segment model.
@@ 6 Conclusion
Adding external high-resolution RGB cameras to the event setup could further improve the sharpness of the rendered novel views. Future work could also investigate architectures that achieve similar quality at smaller cost (e.g. a variant of 3DGS [17]).
@@ 0.5 FPS 1 FPS 5 FPS 10 FPS 50 FPS 100 FPS Ground Truth
Figure 11: Additional synthetic-data ablation on the number of views used for training. With more input views, the quality gets better, albeit hitting diminishing returns at 24 views.
@@ Appendix I Accumulation Stability
To see that noise events can destroy all information about the true brightness level, let us consider a pixel of constant brightness, that nevertheless reports noise events with polarities $p_{i}\in\{-1,+1\}$. For the sake of simplicity, we assume that the $p_{i}$ are independent and identically distributed random variables with fixed, but arbitrary probabilities ${{q}^{+}}:=\mathbb{P}(p_{i}=+1)$ and ${{q}^{-}}:=\mathbb{P}(p_{i}=-1)$. The accumulated polarity after $n$ such events can then be expressed as the random variable
@@
The smaller one chooses $b$, the smaller $\lim\limits_{n\rightarrow\infty}\mathbb{V}(\hat{E}_{n})$ becomes and the faster it converges. However, since decay affects not only noise events but also legitimate ones, we have to explain why it does not distort the actual signal too much. To see this, consider a single pixel: As long as the foreground is not moving through this pixel, it shows a constant level of background brightness. All events it emits are noise, and can safely be suppressed by decay (because the accumulated polarity is supposed to be zero). When the foreground enters the pixel, it will trigger a number of legitimate events. As long as these are recent, the decay will not suppress them too much, so $\hat{E}$ is approximately at foreground brightness level. Assuming that $b$ is chosen suitably, motion is usually fast enough that the object will leave the pixel again before decay “undoes” the entry events completely. Leaving the pixel again triggers a set of legitimate events, that are first sufficiently recent to not be suppressed by damping. As they move further into the past, so do the entry events, so $\hat{E}$ will then correctly approximate the background brightness level again. Of course, decay does still negatively impact the legitimate events (especially when motion is occasionally slower than what $b$ was tuned for, such that the foreground is inside the pixel for longer durations), but the underlying MLP, supervised by all our losses, can compensate for that sufficiently. The ablations in Tabs. 3 and 4 show that SSIM scores, LPIPS and visual results are indeed improved by our decay technique.

### v4

@@ Abstract
Volumetric reconstruction of dynamic scenes is an important problem in computer vision. It is especially challenging in poor lighting and with fast motion. This is partly due to the limitations of RGB cameras: To capture frames under low lighting, the exposure time needs to be increased, which leads to more motion blur. In contrast, event cameras, which record changes in pixel brightness asynchronously, are much less dependent on lighting, making them more suitable for recording fast motion. We hence propose the first method to spatiotemporally reconstruct a scene from sparse multi-view event streams and sparse RGB frames. We train a sequence of cross-faded time-conditioned NeRF models, one per short recording segment. The individual segments are supervised with a set of event- and RGB-based losses and sparse-view regularisation. We assemble a real-world multi-view camera rig with six static event cameras around the object and record a benchmark multi-view event stream dataset of challenging motions. Our work outperforms RGB-based baselines, producing state-of-the-art results, and opens up the topic of multi-view event-based reconstruction as a new path for fast scene capture beyond RGB cameras.
@@ 1 Introduction
In this work, we explore the use of multi-view event data for event-based 4D reconstruction of non-rigid motion. Since a multi-view event camera setup is, of course, more complex than a monocular setup, we believe that investigating this uncharted territory can prove a helpful guideline in deciding whether the added cost is worth the gains in quality.
@@ 2 Related Works
Novel view synthesis has seen recent progress [26, 51, 23, 27, 17], with methods based on Neural Radiance Fields (NeRF) [27] being especially successful. NeRF has been adapted for dynamic scene reconstruction, e.g. head avatars [45], full-body avatars [41, 22, 12], hands [32] and other domains. Other techniques target general scene reconstruction, which can be divided into monocular [43, 40, 52, 20] and multi-view methods [9, 4, 50, 42, 58, 54]. While the monocular methods can reconstruct the input video with high fidelity, the ambiguities of monocular input make it hard for them to learn the correct geometry, greatly limiting how far novel views can deviate from the training viewpoints. Moreover, many monocular methods greatly depend on the camera moving faster than the object, constraining the scenes they can reconstruct [11]. As we capture fast motion, moving the camera even faster is not a viable option. Our method uses a fixed camera multi-view setup. Multi-view methods generalise better to novel views, but are still limited by their RGB input, which is prone to motion blur and noise, especially with fast motion or low lighting conditions. Instead, we use event cameras, possessing both high temporal resolution and high dynamic range. To train NeRF models, one typically needs dozens of ground-truth RGB views, but in our setting, we only have six views available. Follow-up works of NeRF aim at enabling few-view reconstruction, using extensive regularisation [36], pre-training [62, 33], semantic consistency [15], or depth priors [34, 47]. Our approach (Sec. 4.3) uses a combination of regularisers and volume clipping, without a need for pretraining or additional priors, and is applicable to general scenes.
@@
EvDNeRF [1] allows to synthesise events from novel viewpoints using multi-view event streams through the deformation of a canonical volume. It does not model the appearance, needs as many as 18 views, and the deformation-based approach significantly limits the scenes and motions that can be reconstructed. In contrast, our Dynamic EventNeRF allows to reconstruct general dynamic scenes directly in RGB space while using as few as three views.
@@ 3.1 Event Camera Model and Deblurring
Instead of capturing dense RGB frames at regular intervals, event cameras operate asynchronously per individual pixel: When a certain pixel gets brighter or darker by a certain threshold, it emits an event, which is a tuple $(t,x,y,p)$ of its timestamp $t$, pixel position $x,y$, and polarity $p\in\{-1,+1\}$. The event can be interpreted as the assertion
@@
where $\log I_{x,y}(\tau)$ is the intensity of pixel $(x,y)$ at time $\tau$ and $t^{\prime}<t$ is the timestamp of the previous event in that pixel. We call $C_{\mathrm{+1}},C_{\mathrm{-1}}>0$ the event generation thresholds. Event-based Single Integral (ESI) [49] connects the difference between logarithmic intensities on its left side and the accumulated events on the right side by summing Eq. 1 over time. ESI can be further extended to Event-based Double Integral (EDI) [37] by averaging the intensity over the exposure period, to model motion blur caused by the traditional camera shutter. Given events $\mathbf{E}$ and the blurry image $\mathbf{B}$, one can solve a system of equations generated by EDI for the unknown instantaneous intensity image $\mathbf{I}$, thus deblurring $\mathbf{B}$. This way, EDI can be used for deblurring the RGB stream through events. To synthesise arbitrary instantaneous frames $\mathbf{I}(t)$, one could accumulate events on top of the closest recovered deblurred frame $\mathbf{I}(t_{0})$ with ESI. To improve the speed of deblurring and frame synthesis, Lin et al. propose Fast EDI [21], which uses the same event camera model and a special data structure to accelerate the queries.
@@ 4.2 Multi-Segment Model
Using a single MLP to represent longer sequences is challenging due to the capacity-performance trade-off and the ghosting artefacts caused by using a shared network. Therefore, we propose to split the full sequence into short segments and train one model per such sequence. To eliminate flickering caused by changing from one model to the next, we overlap the segments by $10\%$ at the starts and at the ends. In the overlapped parts, we render both models as $\mathbf{\hat{C}}_{i}(t)$ and $\mathbf{\hat{C}}_{i+1}(t)$ and cross-fade the predictions:
@@ 4.3 Supervision for One Segment
For sparse-view reconstruction, we use the following combination of regularisation techniques: First, we want to prevent the model from learning artefacts in the regions where only one view observes them. These areas are near and far away from the camera, but still in its frustum. Suppose the views are distributed 360°360°360  ° around the object at roughly similar height: The intersection of all frustums can be approximated by a cylinder of radius $r$ that extends vertically from $y_{\mathrm{min}}$ to $y_{\mathrm{max}}$. We can then clip the reconstruction volume by setting $\sigma(x,y,z):=0$ wherever $x^{2}+z^{2}>r$, or $y>y_{\textrm{max}}$ or $y<y_{\textrm{min}}$. We also modify the ray depth sampling so that the samples are only chosen inside the cylinder. Following EventNeRF, we randomly offset the ray direction within a pixel to improve generalisation to novel views. This cylinder clipping and the improved sampling procedure result in a strong prior significantly reducing the ambiguity of the sparse-view reconstruction.
@@
Learning high-frequency details from the start is not optimal, especially in the sparse-view setting. Thus, we implement coarse-to-fine training by applying positional encoding annealing [39] both across spatial frequencies and time. Annealing improves temporal and spatial smoothness as the method starts with a coarse representation, slowly progressing towards a finer representation.
@@ 4.4 Event Accumulation
$E_{x,y}(t_{0},t_{1})=\sum_{i=i_{\mathrm{start}}}^{i_{\mathrm{end}}}\delta_{x_{i},y_{i}}^{x,y}p_{i}C_{p_{i}}b^{i_{\mathrm{end}}-i_{\mathrm{start}}-i},~{}b=0.93,$ (11)
@@
In Eq. 6, we defined $\mathcal{L}_{\mathrm{event}}$, which uses $E(t_{0},t_{1})$ for random $t_{0}$ and $t_{1}$ each training iteration. As event streams could contain millions of events per second, implementing the accumulation procedure directly by processing each event on every query would be too slow to run every training iteration. To overcome this issue, we use an approach inspired by Lin et al. [21] to accelerate the queries: We treat all pixels independently and store all intermediate accumulation results in arrays $A^{x_{i},y_{i}}_{j}=E(0,t_{i})_{x_{i},y_{i}}$, and $T^{x_{i},y_{i}}_{j}=t_{i}$, where $j$ is the per-pixel event counter. Then, once we need to compute $E(t_{0},t_{1})$, we split it into two queries:
@@ 5.5 Ablation and Design Choice Study
We conduct an ablation and design choice study (Tabs. 2, 5 and 4; per-scene in Tab. 4 in the appendix) using both synthetic and real data. In particular, we compare different options for the core model: MLP (“Full Model”), NGP [31], TensoRF-CP [7], and HexPlane [4]. NGP fails to handle the sparse-view setting. TensoRF-CP and HexPlane are both grid-based: Individual timestamps are modelled separately, so that information from one timestamp cannot propagate to other timestamps, leading to blurriness and artefacts. On the other hand, our full model with temporally-conditioned MLP explicitly controls how much information is shared between timestamps through positional encoding. Supervising one timestamp also affects others in its vicinity, resulting in better use of the training data. Disabling $\mathcal{L}_{\text{event}}$, $\mathcal{L}_{\text{acc}}$, or $\mathcal{L}_{\text{RGB}}$ leads to increased blurriness because disabling either of them will reduce the amount of supervision: Without $\mathcal{L}_{\text{event}}$, there is no way to correct relative changes between predictions through events, without $\mathcal{L}_{\text{acc}}$ there is no way to propagate sharp RGB information and static parts of the scene. And without $\mathcal{L}_{\text{RGB}}$, the model is not encouraged to predict the reference RGB frames correctly. Disabling both $\mathcal{L}_{\text{RGB}}$ and $\mathcal{L}_{\text{acc}}$ removes the entire RGB supervision. As the model only uses the events, there can be an exposure ambiguity, leading to lower metrics. Disabling both $\mathcal{L}_{\text{RGB}}$ and $\mathcal{L}_{\text{event}}$ leaves only $\mathcal{L}_{\text{acc}}$. Despite using both RGB and events, this baseline performs worse than the full model, presumably due to accumulation artefacts. Cylinder clipping prevents the model from creating geometry in regions that are only supervised by one view. If disabled, the model produces numerous artefacts in the novel views; in the case of real data specifically, it even fails to converge. $\mathcal{L}_{\text{sparsity}}$ forces the model to reduce the number of opaque pixels in the reconstruction. Removing it leads to semi-transparent floating artefacts in the reconstructed volume. Using only one single model for the entire sequence duration (as opposed to dividing the sequence into shorter chunks) results in a blurrier reconstruction. In this case, artefacts produced by reconstructing one of the segments can propagate to other segments, which is not the case in the multi-segment model.
@@ 6 Conclusion
Adding external high-resolution RGB cameras to the event setup could further improve the sharpness of the rendered novel views. Future work could also investigate architectures that achieve similar quality at a smaller cost (e.g. a variant of 3DGS [17]).
@@ 0.5 FPS 1 FPS 5 FPS 10 FPS 50 FPS 100 FPS Ground Truth
Figure 11: Additional synthetic-data ablation on the number of views used for training. With more input views, the visual quality improves, albeit hitting diminishing returns at 24 views.
@@ Appendix I Accumulation Stability
To see that noise events can destroy all information about the true brightness level, let us consider a pixel of constant brightness that, nevertheless, reports noise events with polarities $p_{i}\in\{-1,+1\}$. For the sake of simplicity, we assume that the $p_{i}$ are independent and identically distributed random variables with fixed, but arbitrary probabilities ${{q}^{+}}:=\mathbb{P}(p_{i}=+1)$ and ${{q}^{-}}:=\mathbb{P}(p_{i}=-1)$. The accumulated polarity after $n$ such events can then be expressed as the random variable
@@
The smaller one chooses $b$, the smaller $\lim\limits_{n\rightarrow\infty}\mathbb{V}(\hat{E}_{n})$ becomes and the faster it converges. However, since decay affects not only noise events but also legitimate ones, we have to explain why it does not distort the actual signal too much. To see this, consider a single pixel: As long as the foreground is not moving through this pixel, it shows a constant level of background brightness. All events it emits are noise, and can safely be suppressed by decay (because the accumulated polarity is supposed to be zero). When the foreground enters the pixel, it will trigger a number of legitimate events. As long as these are recent, the decay will not suppress them too much, so $\hat{E}$ is approximately at the foreground brightness level. Assuming that $b$ is chosen suitably, motion is usually fast enough that the object will leave the pixel again before decay “undoes” the entry events completely. Leaving the pixel again triggers a set of legitimate events that are first sufficiently recent not to be suppressed by damping. As they move further into the past, so do the entry events, so $\hat{E}$ will then correctly approximate the background brightness level again. Of course, decay does still negatively impact the legitimate events (especially when motion is occasionally slower than what $b$ was tuned for, such that the foreground is inside the pixel for longer durations), but the underlying MLP, supervised by all our losses, can compensate for that sufficiently. The ablations in Tabs. 3 and 4 show that SSIM scores, LPIPS and visual results are indeed improved by our decay technique.